OpenStudy (anonymous):

Points P(-2,4,3) and Q(-3,0,3) lie on line L. Does a point exist where line L intersects the XZ-Plane? If it does what point.

OpenStudy (anonymous):

So we have the vector PQ = <-1,-4,0>

OpenStudy (anonymous):

And a point is on the xz plane if it has 0 as a z coordinate like so: (x, y, 0)

OpenStudy (amistre64):

P(-2,4,3) and Q(-3,0,3) +2-4-3 +2-4-3 ---------------------- 0 0 0 <-1,-4, 0> we need any scalar of this vector soo: <-t,-4t, 0> and lets apply it to one of the points given

OpenStudy (amistre64):

x=-2-t y= 4-4t z= 3 is our line

OpenStudy (anonymous):

line equation : (-2,4,3) + (-1,-4,0)t now we need to find t such that (-2-t,4-4t,3) 4-4t = 0 (XZ PLANE) so t =1 the point is (-2-1,0,3) -> (-3,0,3)

OpenStudy (amistre64):

in order to cut thru the xz plane; y=0

OpenStudy (amistre64):

y = 0 what t = 1 x=-2-1 = -3 y= 4-4(1)= 0 z= 3 = 3

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

I am struggling with this stuff, its mad complicated, you guys have any tips?

OpenStudy (anonymous):

*or girls

OpenStudy (amistre64):

.... i have tips, no girls tho that i can give you

OpenStudy (anonymous):

is it high school ? (im wondering .. where i live we dont study this in high school)

OpenStudy (anonymous):

Nope, community college.

OpenStudy (amistre64):

the biggest thing is to draw a picture so that you can see it